# Johannes Matzat        Aiko Schmeißer

### v4

@@ 2.1 Unionizing through NLRB Elections
Since 1935, the National Labor Relations Act (NLRA) gives most private-sector workers in the U.S. the right to organize in unions and take collective action, such as bargaining and strikes. Collective bargaining between unions and employers takes place at the establishment level. Traditionally, workers unionize through a secret ballot election at their establishment that is administered by the National Labor Relations Board (NLRB). [^7] While union elections are the primary means by which private-sector workers gain union representation, there are alternative procedures for unionization. First, employers may voluntarily recognize unions without an election through neutrality agreements and “card checks”. These cases are less common, however, since employers generally oppose union organization (Schmitt and Zipperer, 2009). Second, some workers’ bargaining rights are not regulated by the NLRA. For example, the Railway Labor Act determines bargaining rights of airline and railroad workers and several federal, state, and local laws regulate the organization of public-sector employees. The unionization procedure involves three main steps: a petition drive, an election, and certification.
The organizing drive can be initiated either by the workers at an establishment or by a union organization. [^8] See Prummer and Squintani (2023) for an organizational theory on how unionization success can vary between grassroots movements and top-down approaches by established unions. The initiator first needs to gather the signatures of at least 30% of workers in the proposed bargaining unit who thereby express a desire for unionization. With these signatures, an election petition is filed to the NLRB. The NLRB decides whether to accept the petition by ascertaining whether workers in the proposed bargaining unit share common interests that can be adequately represented by the union. If the petition is accepted, the NLRB schedules a secret ballot election, which usually takes place in the workplace. The union wins the election if it obtains a strict majority of the votes cast. In case of union victory, the NLRB certifies the union as the sole authorized representative of employees in the bargaining unit.
Union certification requires the employer to bargain “in good faith” with the union. This bargaining generally aims at concluding a first contract between the union and the employer. However, as there is no legal obligation to reach such an agreement, only about 55% of certifications yield a first contract within two years of the election (Ferguson, 2008). When both parties cannot reach a first agreement (or when subsequently they are disputing over the terms and conditions of the first contract), they can consult a neutral third party to resolve disputes via mediation or arbitration. After one year has passed since certification, employees can also decide to hold a decertification election to vote out the union.
@@
Money plays a dominant role in U.S. politics. Monetary resources are viewed as essential for political candidates in order to take part and be successful in the political process. There is indeed increasing evidence that campaign funding can influence who runs for and who wins elections (e.g., Avis et al., 2022; Barber, 2016b; Battaglini et al., 2024; Broberg et al., 2023). The large majority of campaign contributions in the U.S. originate from individual donors. The share of contributions to candidates that are donated by individuals has increased from 54% in 2000 to 77% in the 2020 congressional election cycle (FEC, 2022a). While political spending is certainly concentrated among the wealthy (Bonica and Rosenthal, 2018; Hill and Huber, 2017), it is a prevalent form of political participation for a substantial share of the U.S. electorate. Bouton et al. (2024) estimate that 12.7% of the adult U.S. citizen population have made at least one campaign contribution between 2006 and 2020.
@@
Unlike corporations, which are prohibited by U.S. federal law to support candidates directly out of treasury funds, individual donors are allowed to make direct contributions to political candidates. [^10] To make campaign donations, companies must set up a PAC, which may only solicit contributions from the firm’s employees. The PAC can in turn donate directly to political candidates or other recipients. There are, however, restrictions to the maximum amount that an individual can donate to a candidate. The limit varies by recipient type and election cycle. For the 2020 federal elections, for example, individuals were allowed to donate at most 2,800 USD to a single candidate and 5,000 USD to a PAC. Recipients are obligated to itemize all individual contributions greater than 200 USD and report the donor’s identifying information along with the amount and date of the contribution. Donations smaller than 200 USD are not required to be itemized but are included in the total amount that the recipient reports to the Federal Election Commission. [^11] Over the last years, increasing shares of donations are channeled through party’s online fundraising platforms, such as Democrats’ ActBlue (created in 2004) and Republicans’ WinRed (created in 2019). Importantly, these platforms do not have a minimum reporting threshold and thus also have to itemize contributions below 200 USD. As a result, in recent years we can increasingly observe information on small donations (Bouton et al., 2024).
@@
The literature differentiates two broad motivations for why individuals contribute to political candidates. First, contributions can be seen as consumption goods that give individuals consumption value from participating in politics and sponsoring candidates that are ideologically close to their own political position (Ansolabehere et al., 2003). Second, donors may view contributions as investment goods that can buy access to politicians and benefit their own material interests. There is extant evidence that individuals’ donations are ideologically motivated. Individual donors self-report that candidate ideology has great importance when deciding to whom to give (Barber, 2016a). Moreover, in comparison to access-seeking PACs, who prefer donating to moderate candidates, individuals tend to support more ideologically extreme candidates (Barber, 2016b; Stone and Simas, 2010). In merged survey-administrative data, contribution-based ideology measures also predict policy preferences of donors, even of donors from the same party (Bonica, 2018). Most related to our study, Stuckatz (2022) analyzes the role of the workplace in shaping employees’ contribution behavior by examining alignment in donations of employees and their employer PACs. He finds that 13% of donations from rank-and-file workers and 21% of donations from executives go to candidates endorsed by their employer PACs, suggesting that only a small fraction of employees’ donations may be influenced by corporate strategic objectives. Teso (forthcoming) shows that a business leader’s likelihood of donating to a Member of Congress increases when the politician becomes assigned to a committee that is policy-relevant to the business leader’s company. In quantitative terms, his estimates imply that less than 2% of corporate leaders’ donations to Congress members are driven by influence-seeking motives. Moreover, Bonica (2016) finds that donations from corporate board members are ideologically quite diverse, both across and within companies. Compared to corporate PACs, business leaders also tend to support more non-incumbent candidates and less powerful legislators. In sum, the evidence supports ideological preferences as the main driver of individuals’ campaign donations.
@@ 3 Data
Previous studies have been unable to assess the political impact of unions at the establishment level due to a lack of matched employer-employee data for political outcomes. Campaign contribution data are uniquely suited to overcome this constraint. To ensure transparency in politicians’ campaign funds, contributors are required to disclose their name, employer, address, and occupation. The employer and location information allows us to link donors to the union election results of their employers. We are not aware of any other large-scale data on political behavior with employer information in the U.S. that would allow this link. Furthermore, we can use the occupation information to study the political effects of unionization not only on directly affected non-managerial workers but also on potentially indirectly affected managers and supervisors. In the following, we describe how we construct a new establishment-level dataset that links union elections to campaign contributions from employees.
@@ 3.1 Union Elections
We start with a comprehensive dataset on the universe of U.S. union representation elections between 1961 and 2018. Specifically, we combine data collected by Farber (2016) with public data from NLRB election reports. [^12] We obtain the dataset originally assembled by Farber (2016) from the replication package of Knepper (2020). The data contain information on elections held between 1961 and 2009. For elections between 2010 and 2018, we retrieve data from NLRB election reports (link). Each data point represents a union election at a single establishment and contains vote counts for and against unionization, the dates of the petition filing and of the actual election, as well as the name of the union organization. Moreover, it includes the establishment’s name and address, which we exploit to match campaign contributions.
@@
Sample restrictions. Before matching elections to campaign contributions, we impose several sample restrictions. [^13] See Appendix B.1 for details on the union election data and the sample restrictions we impose. First, we only consider elections held between 1985 and 2010. Given that our contribution data cover the years 1979-2016, this allows us to observe trends in contributions for three political election cycles (six years) before and after each union election. Second, we follow Frandsen (2021) and restrict the sample to union elections where at least 20 votes were cast. This restriction ensures that winning establishments are affected by a non-trivial rise in union representation. Moreover, it helps to exclude small establishments, which are more likely to have come into existence recently and have a lower probability of survival over our period of analysis. Third, following Knepper (2020) and Wang and Young (2023), we only keep the first union election in each establishment. [^14] In the election data, we identify an establishment as a unique address or a unique combination of the standardized firm name and commuting zone. For a firm that has multiple establishments within the same commuting zone, we thus only consider the first election among these establishments. Excluding non-inaugural elections avoids having multiple observations for the same establishment with reversed treatment status over time, and helps alleviate election manipulation issues if managers or unions learn how to apply manipulation tactics in repeat elections. Our estimates should thus be interpreted as the effects of winning the first union election. [^15] This does not perfectly correspond to the effect of union representation in all post-election periods for two reasons. First, establishments may lose representation after a decertification election. Wang and Young (2023) show that 5-10% of establishments that win a first union election hold a decertification election within 5 years. Second, establishments, after losing the first election, can hold another successful election in subsequent years. According to DiNardo and Lee (2004), this is the case for around 10% of lost first elections. By focusing on the effect of winning the first election, we thus accept an attenuation of our estimates relative to the true effects of union representation over all post-election periods. These restrictions leave us with a sample of 28,823 union elections, which we seek to match to the campaign contribution data.
@@ 3.2 Campaign Contributions
To measure the political mobilization and ideology of employees, we use the Database on Ideology, Money in Politics and Elections (DIME) compiled by Bonica (2019). [^16] Other papers have used these data to study, among others, the political consequences of import competition (Autor et al., 2020) and immigration (Dreher et al., 2023). DIME provides transaction-level data on campaign contributions registered with the Federal Election Commission and other state and local election commissions. We exploit all campaign contributions from individuals to candidates running for office at the federal and local level (specifically the President, House of Representatives, Senate, Governor, and upper and lower chambers of state legislature), as well as to all PACs (including party and interest group PACs). The dataset covers the 1979-2016 period and includes the amount and exact date of the donation, as well as identifying information on the donor and recipient. [^17] Accurate reporting of the donor information (name, employer, address, occupation) is enforced by the Federal Election Commission through regular audits, as well as fines and further legal action in case of non-compliance. See FEC (2022b) for enforcement statistics.
@@
Bonica (2019) deploys identity resolution techniques to assign unique identifiers to each donor. The identifiers allow us to track donors’ contributions over time, which we exploit to study whether establishment-level effects are driven by compositional changes from leaving and newly hired employees or by individual-level effects on incumbent employees. Further, the DIME includes measures for the political ideology of recipients and donors, so-called campaign finance (CF) scores, which are derived by Bonica (2014) from solving a spatial model of contributions. The model formalizes the idea that donors contribute more to candidates with a similar ideological position and estimates ideal points of both recipients and donors along a typical liberal-conservative scale. Using the ideology scores, we can go beyond previous papers that only relate unions to Democratic versus Republican party affiliation and study how unionization affects ideological contribution patterns for candidates within the same party.
@@
Matching algorithm. We link the campaign contributions to the employing establishments with union elections by combining a spatial match with a fuzzy match of firm names. We start by restricting potential matches to the same local labor market using 1990 commuting zones. 92% of the population live and work in the same local labor market, making it very likely that a donor in our sample works at an establishment in the same local labor market (Fowler and Jensen, 2020). The restriction substantially reduces the computational requirements for the fuzzy match and ensures that for multi-establishment firms we do not incorrectly match employees to establishments of the same firm in other locations. [^18] We accept measurement error from assigning donors to the wrong establishment if a firm has several establishments within a commuting zone. However, within-firm interactions may generate spillover effects across establishments. The results of Knepper (2020), for example, imply large spillovers in the effects of unionization on firm-level employee compensation. To match the employer name in the contribution data to the establishment name in the union election data, we use an automated record-linkage program introduced by Blasnik (2010) and Wasi and Flaaen (2015). The linkage process first standardizes employer names and then calculates bigram scores for the similarity of each string pair. Lastly, we manually review all matches with a score above a minimum threshold. [^19] See Appendix B.2 for details on the matching process.
@@
To arrive at an establishment-level panel of employee contributions, we sum up all matched contributions within an establishment and two-year election cycle. Our period of analysis covers three cycles before to three cycles after each union election. In our baseline estimation sample, we only include establishments for which we observe at least one matched contribution over this period in order to reduce bias from false negative matches. [^20] Note that we generally took a conservative approach in the manual review of matches and rejected potentially true matches if firm names in the union election data were misspelled or too generic to infer a unique firm per commuting zone (e.g. ”community health center”, ”general construction”, ”support services corporation”). Including these establishments and assigning them zero amounts in the estimation sample, would lead to an attenuation bias in our estimates. Appendix Table A.1 compares the characteristics of establishments with and without at least one matched contribution. Elections in our matched sample involve more voters, i.e., are likely to be larger, and tend to be held in more recent years as contribution numbers have sharply increased over time. At the same time, the matching does not strongly affect the selection of union elections in terms of voting outcome and industry composition. In Appendix Table A.2, we verify that unionization does not affect the likelihood of observing any contribution after the union election, suggesting that the sample selection criterion is not related to our treatment of interest. Moreover, we will discuss the sensitivity of our main results to alternative sample restrictions. For our baseline sample, we are left with 6,063 matched establishments (and 42,441 establishment-cycle observations). As Table 1 reports, the sample is built from 357,436 matched contributions that amount to 105.8 million USD spent by 46,719 different donors to 9,942 different recipients.
@@
Classification of occupations. In order to differentiate between workers eligible for unionization and their managers and supervisors who are always excluded from the bargaining unit, we classify self-reported occupations of donors. Here, we only briefly describe the classification procedure and provide more details in Appendix B.3. We start by mapping the free-text occupation descriptions in the DIME to the 6-digit Standard Occupation Classification (SOC). For this, we combine an ensemble classifier called SOCcer (Russ et al., 2016), sub- and fuzzy string matching to an extensive crosswalk of laymen’s occupation titles from O*NET, as well as manual reviews of the most common occupation titles. Appendix Figure A.1 shows the occupation distribution for the classified donations. While the largest share (44%) is given by donors in management occupations, we also see substantial shares of contributions originating from lower-tier white-collar occupations such as healthcare, education, culture and sports, or financial operations workers. Blue-collar occupations, in contrast, account for small shares of the overall number of contributions, which is not surprising given that wealth is a strong predictor of political donating.
@@
With the classified SOC codes at hand, we categorize donors into managers and supervisors versus non-managerial workers. We identify managers and supervisors by using all contributions from “Management Occupations” (SOC group 11) and adding all occupations that involve a significant amount of supervising following the NLRA definition of supervisor tasks and leveraging occupational task descriptions from O*NET. Non-managerial workers are then defined as all remaining donors to whom we were able to assign a SOC code. The occupational composition in our final sample of candidate contributions looks as follows: 42% of contributions originate from managers and supervisors (hereafter only termed “managers”), 30% from non-managerial workers (hereafter only termed “workers”), and for 28% we are unable to obtain an occupational classification. Due to the non-negligible share of unclassified occupations, we report results not only separately for managers and workers, but also for all employees together (including those without a classification). [^21] In Appendix B.3 we also provide evidence that the likelihood of having a missing occupation classification is not affected by unionization and therefore unlikely to drive our results.

### v5

@@ 2.1 Unionizing through NLRB Elections
Since 1935, the National Labor Relations Act (NLRA) gives most private-sector workers in the U.S. the right to organize in unions and take collective action, such as bargaining and strikes. Collective bargaining between unions and employers takes place at the establishment level. Traditionally, workers unionize through a secret ballot election at their establishment that is administered by the National Labor Relations Board (NLRB). [^10] While union elections are the primary means by which private-sector workers gain union representation, there are alternative procedures for unionization. First, employers may voluntarily recognize unions without an election through neutrality agreements and “card checks”. These cases are less common, however, since employers generally oppose union organization (Schmitt2009). Second, some workers’ bargaining rights are not regulated by the NLRA. For example, the Railway Labor Act determines bargaining rights of airline and railroad workers and several federal, state, and local laws regulate the organization of public-sector employees. The unionization procedure involves three main steps: a petition drive, an election, and certification.
The organizing drive can be initiated either by the workers at an establishment or by a union organization. The initiator first needs to gather the signatures of at least 30% of workers in the proposed bargaining unit who thereby express a desire for unionization. With these signatures, an election petition is filed to the NLRB. The NLRB decides whether to accept the petition by ascertaining whether workers in the proposed bargaining unit share common interests that can be adequately represented by the union. If the petition is accepted, the NLRB schedules a secret ballot election, which usually takes place in the workplace. The union wins the election if it obtains a strict majority of the votes cast. In case of union victory, the NLRB certifies the union as the sole authorized representative of employees in the bargaining unit.
Union certification requires the employer to bargain “in good faith” with the union. This bargaining generally aims at concluding a first contract between the union and the employer. However, as there is no legal obligation to reach such an agreement, only about 55% of certifications yield a first contract within two years of the election (Ferguson2008). When both parties cannot reach a first agreement (or when subsequently they are disputing over the terms and conditions of the first contract), they can consult a neutral third party to resolve disputes via mediation or arbitration. After one year has passed since certification, employees can also decide to hold a decertification election to vote out the union.
@@
Money plays a dominant role in U.S. politics. Monetary resources are viewed as essential for political candidates in order to take part and be successful in the political process. There is indeed increasing evidence that campaign funding can influence who runs for and who wins elections (e.g., Avis2022; Barber2016a; Battaglini2024; Broberg2023). The large majority of campaign contributions in the U.S. originate from individual donors. The share of contributions to candidates that are donated by individuals has increased from 54% in 2000 to 77% in the 2020 congressional election cycle (FEC2022a). While political spending is certainly concentrated among the wealthy (Bonica2018a; Hill2017), it is a prevalent form of political participation for a substantial share of the U.S. electorate. Bouton2021 estimate that 12.7% of the adult U.S. citizen population have made at least one campaign contribution between 2006 and 2020.
@@
Unlike corporations, which are prohibited from supporting candidates directly out of treasury funds, individual donors are allowed to make direct contributions to political candidates. [^12] To make campaign donations, companies must set up a PAC, which may only solicit contributions from the firm’s employees. The PAC can in turn donate directly to political candidates or other recipients. There are, however, restrictions to the maximum amount that an individual can donate to a candidate. The limit varies by recipient type and election cycle. For the 2024 federal elections, for example, individuals were allowed to donate at most 3,300 USD to a single candidate and 5,000 USD to a PAC. Recipients are obligated to itemize all individual contributions greater than 200 USD and report the donor’s identifying information along with the amount and date of the contribution. Donations smaller than 200 USD are not required to be itemized but are included in the total amount that the recipient reports to the Federal Election Commission. [^13] Over the last years, increasing shares of donations are channeled through party’s online fundraising platforms, such as Democrats’ ActBlue (created in 2004) and Republicans’ WinRed (created in 2019). Importantly, these platforms do not have a minimum reporting threshold and thus also have to itemize contributions below 200 USD. As a result, in recent years we can increasingly observe information on small donations (Bouton2021).
@@
The literature differentiates two primary motivations for why individuals contribute to political candidates. First, contributions may serve as consumption goods, providing intrinsic value to donors who participate in politics and support candidates ideologically close to their own political position (Ansolabehere2003). Second, donors may view contributions as investment goods that can buy access to politicians and benefit their own material interests. There is extant evidence that individuals’ donations are ideologically motivated. Donors report that candidate ideology has great importance when deciding to whom to give (Barber2016). Moreover, unlike access-seeking PACs, who prefer donating to moderate candidates, individuals tend to support more ideologically extreme candidates (Barber2016a; Stone2010). In merged survey-administrative data, contribution-based ideology measures also predict policy preferences of donors, even of donors from the same party (Bonica2018). That said, recent studies have also explored the role of influence-seeking motives in individual campaign contributions. Teso2023 shows that a business leader’s likelihood of donating to a Member of Congress increases when the politician is assigned to a committee that is policy-relevant to the business leader’s company. However, these effects are quantitatively limited: the estimates imply that less than 2% of corporate leaders’ donations to Congress members are driven by influence-seeking motives. [^14] Further speaking against a strong role of influence-seeking motives, Bonica2016 finds that donations of corporate board members are ideologically quite diverse, both across and within companies. Compared to corporate PACs, business leaders also tend to support more non-incumbent candidates and less powerful legislators. Stuckatz2022 analyzes how the workplace shapes employees’ contribution behavior by examining alignment between donations of employees and those of their employer PACs. While the paper documents positive associations, the results indicate that only 13% of donations from rank-and-file workers and 21% of donations from executives go to candidates endorsed by their employer PACs, suggesting that a limited share of employees’ donations can be influenced by corporate strategic objectives. Taken together, the evidence supports the interpretation that ideological preferences, rather than strategic influence-seeking, are the predominant motivation behind individuals’ campaign contributions.
@@ 3 Data
Previous studies have been unable to assess the political impact of unions at the establishment level due to a lack of matched employer-employee data for political outcomes. Campaign contribution data are uniquely suited to overcome this constraint. To ensure transparency in politicians’ campaign funds, contributors are required to disclose their name, employer, address, and occupation. The employer and location information allows us to link donors to the union election results of their employers. Furthermore, we can use the occupation information to study the political effects of unionization not only on directly affected non-managerial workers but also on potentially indirectly affected managers and supervisors. In the following, we describe how we construct a new establishment-level dataset that links union elections to campaign contributions from employees.
@@ 3.1 Union Elections
We start with a comprehensive dataset on the universe of U.S. union representation elections between 1961 and 2018. Specifically, we combine data collected by Farber2016 with public data from NLRB election reports. [^15] We obtain the dataset originally compiled by Farber2016 from the replication package of Knepper2020, covering elections from 1961 to 2009, and add data from NLRB election reports for the years 2010 to 2018. See Appendix B.1 for details on the union election data, including the sample restrictions we impose. Each data point represents a union election at a single establishment and contains vote counts for and against unionization, the dates of the petition filing and of the actual election, as well as the name of the union organization. Moreover, it includes the establishment’s name and address, which we exploit to match campaign contributions.
@@
Sample restrictions. Before matching elections to campaign contributions, we impose several sample restrictions. First, we only consider elections held between 1985 and 2010. Given that our contribution data cover the years 1979-2016, this allows us to observe trends in contributions for three political election cycles (six years) before and after each union election. Second, we follow Frandsen2020 and restrict the sample to union elections where at least 20 votes were cast. This restriction ensures that winning establishments are affected by a non-trivial rise in union representation. Moreover, it helps to exclude small establishments, which are more likely to have come into existence recently and have a lower probability of survival over our period of analysis. Third, following Knepper2020 and Wang2021, we only keep the first union election in each establishment. [^16] In the election data, we identify an establishment as a unique address or a unique combination of the standardized firm name and commuting zone. For a firm that has multiple establishments within the same commuting zone, we thus only consider the first election among these establishments. Excluding non-inaugural elections avoids having multiple observations for the same establishment with reversed treatment status over time, and helps alleviate election manipulation issues if managers or unions learn how to apply manipulation tactics in repeat elections. Our estimates should thus be interpreted as the effects of winning the first union election. [^17] This does not perfectly correspond to the effect of union representation in all post-election periods for two reasons. First, establishments may lose representation after a decertification election. Wang2021 show that 5-10% of establishments that win a first union election hold a decertification election within 5 years. Second, establishments, after losing the first election, can hold another successful election in subsequent years. According to DiNardo2004, this is the case for around 10% of lost first elections. By focusing on the effect of winning the first election, we thus accept an attenuation of our estimates relative to the true effects of union representation over all post-election periods. These restrictions leave us with a sample of 28,823 union elections, which we seek to match to the campaign contribution data.
@@ 3.2 Campaign Contributions
To measure the political mobilization and ideology of employees, we use the Database on Ideology, Money in Politics and Elections (DIME) compiled by Bonica2019. DIME provides transaction-level data on campaign contributions registered with the Federal Election Commission and other state and local election commissions. We exploit all campaign contributions from individuals to candidates running for office at the federal and local level (specifically the President, House of Representatives, Senate, Governor, and upper and lower chambers of state legislature), as well as to all PACs (including party and interest group PACs). The dataset covers the 1979-2016 period and includes the amount and exact date of the donation, as well as identifying information on the donor and recipient. [^18] Accurate reporting of the donor information (name, employer, address, occupation) is enforced by the Federal Election Commission through regular audits, as well as fines and further legal action in case of non-compliance. See FEC2022 for enforcement statistics.
@@
Bonica2019 deploys identity resolution techniques to assign unique identifiers to each donor. The identifiers allow us to track donors’ contributions over time, which we exploit to study whether establishment-level effects are driven by compositional changes from leaving and newly hired employees or by individual-level effects on incumbent employees. Further, the DIME includes measures for the political ideology of recipients and donors, so-called campaign finance (CF) scores, which are derived by Bonica2014 from solving a spatial model of contributions. The model formalizes the idea that donors contribute more to candidates with a similar ideological position and estimates ideal points of both recipients and donors along a typical liberal-conservative scale. Using the ideology scores, we can go beyond previous papers that only relate unions to Democratic versus Republican party affiliation and study how unionization affects ideological contribution patterns for candidates within the same party.
@@
Matching algorithm. We link the campaign contributions to the employing establishments with union elections by combining a spatial match with a fuzzy match of firm names. We start by restricting potential matches to the same local labor market using 1990 commuting zones. 92% of the population live and work in the same local labor market, making it very likely that a donor in our sample works at an establishment in the same local labor market (Fowler2020). The restriction substantially reduces the computational requirements for the fuzzy match and ensures that for multi-establishment firms we do not incorrectly match employees to establishments of the same firm in other locations. [^19] We accept measurement error from assigning donors to the wrong establishment when a firm has several establishments within a commuting zone. Note that within-firm interactions may also generate spillover effects across establishments (Knepper2020; Wang2021). To match the employer name in the contribution data to the establishment name in the union election data, we use the automated record-linkage program of Blasnik2010 and Wasi2015. The linkage process first standardizes employer names and then calculates bigram scores for the similarity of each string pair. Lastly, we manually review all matches with a score above a minimum threshold. [^20] See Appendix B.2 for details on the matching process. We have also experimented with matching employer names using a large language model which yields similar match accuracy in a test dataset.
@@
To arrive at an establishment-level panel of employee contributions, we sum up all matched contributions within an establishment and two-year election cycle. Our period of analysis covers three cycles before to three cycles after each union election. In our baseline estimation sample, we only include establishments for which we observe at least one matched contribution over this period in order to reduce bias from false negative matches. [^21] Note that we generally took a conservative approach in the manual review of matches and rejected potentially true matches if firm names in the union election data were misspelled or too generic to infer a unique firm per commuting zone (e.g. ”community health center”, ”general construction”, ”support services corporation”). Including these establishments and assigning them zero amounts in the estimation sample, would lead to an attenuation bias in our estimates. Appendix Table A.1 compares the characteristics of establishments with and without at least one matched contribution. Elections in our matched sample involve more voters, i.e., are likely to be larger, and tend to be held in more recent years as contribution numbers have sharply increased over time. At the same time, the matching does not strongly affect the selection of union elections in terms of voting outcome and industry composition. In Appendix Table A.2, we verify that unionization does not affect the likelihood of observing any contribution after the union election, suggesting that the sample selection criterion is not related to our treatment of interest. Moreover, we will discuss the sensitivity of our main results to alternative sample restrictions.
For our baseline sample, we are left with 6,063 matched establishments (and 42,441 establishment-cycle observations). As Table 1 reports, the sample is built from 357,436 matched contributions that amount to 105.8 million USD spent by 46,719 different donors to 9,942 different recipients. [^22] Extrapolating from our matched sample, we can estimate the total amount contributed by employees at all unionized establishments. We first calculate the cycle-specific average amount donated per eligible voter in newly unionizing establishments and then multiply it by the number of all workers represented by unions during that cycle. Summing up over all cycles from 1986 to 2010, we estimate that employees at unionized establishments have donated a total of 1.36 billion USD. Contrasting this number with the cost of running successfully for a seat in the House of Representatives – which lies between \$777,000 and \$1.47 million (persily2018campaign) – highlights the sizable potential for political influence that unions could wield by impacting the donations of employees at unionized establishments.
@@
Classification of occupations. In order to differentiate between workers eligible for unionization and their managers and supervisors who are always excluded from the bargaining unit, we classify self-reported occupations of donors. We start by mapping the free-text occupation descriptions in the DIME to the 6-digit Standard Occupation Classification (SOC). For this, we combine an ensemble classifier called SOCcer (Russ2016), sub- and fuzzy string matching to an extensive crosswalk of laymen’s occupation titles from O*NET, as well as manual reviews of the most common occupation titles. [^23] Appendix Figure A.1 shows the occupation distribution for the classified donations. While the largest share (44%) is given by donors in management occupations, we also see substantial shares of contributions originating from lower-tier white-collar occupations such as healthcare, education, culture and sports, or financial operations workers. Blue-collar occupations, in contrast, account for small shares of the overall number of contributions, which is not surprising given that wealth is a strong predictor of political donating. With the classified SOC codes at hand, we categorize donors into managers and supervisors versus non-managerial workers. We identify managers and supervisors by using all contributions from “Management Occupations” (SOC group 11) and adding all occupations that involve a significant amount of supervising following the NLRA definition of supervisor tasks and leveraging occupational task descriptions from O*NET. Non-managerial workers are then defined as all remaining donors to whom we were able to assign a SOC code.
@@
The occupational composition in our final sample of candidate contributions looks as follows: 42% of contributions originate from managers and supervisors (hereafter only termed “managers”), 30% from non-managerial workers (hereafter only termed “workers”), and for 28% we are unable to obtain an occupational classification. Due to the non-negligible share of unclassified occupations, we report results not only separately for managers and workers, but also for all employees together (including those without a classification). Moreover, as a robustness check, we classify workers and managers using information on the median household income in the census tract they live in. [^24] See Appendix B.3 for details on the occupational classification. There, we also provide evidence that the likelihood of having a missing occupation classification is not affected by unionization and therefore unlikely to drive our results.